OpenStudy (anonymous):

5c-4t=8 14+4t=3c solve by elimination

OpenStudy (anonymous):

5c-4t=8 (1) 14+4t=3c (2) Move the variables to one side in equation (2) and the constant on the next

OpenStudy (anonymous):

What equations do you end up with when you do this?

OpenStudy (ajprincess):

\(5c-4t=8-(1)\) \(14+4t=3c\) \(3c-4t=14-(2)\) by subtracting (2) from (1) term containing t will be eliminated nd u will get an equation in terms of c. Then solve it to find c.:)

OpenStudy (anonymous):

subtract two equtions 5c-4t=8 - * 3c-4t=14 2c=-6 c=-3 put in any of the two t=-23/4

OpenStudy (anonymous):

I understand now thank you ^__^.

OpenStudy (ajprincess):

welcome:)